The force of attraction between two point charges at a distance r apart is F. What should be the distance apart in the same medium so that the force becomes F/3?what if force becomes 3F

Grade:12

2 Answers

Adarsh
6 years ago
If the distance between them is increased by 3 times the initial length then force will be F/3.
if the distance is decreased by 1/3 times the initial length then force will be 3F.
Thanks and Regards 
Adarsh
Khimraj
6 years ago
Force of attraction between two charges is inversely proportional to square of distance. So
F1/F2 = r22/r12
3 = r22/r2
So r2 = r*(3)1/2
 Similarly if force become 3F then
r2 = r*(1/3)1/2
Hope it clears.